OpenStudy (anonymous):

I dont know how to do this arctansqrt((− 3 ))

OpenStudy (tkhunny):

Are you sure it's \(\sqrt{-3}\) and not \(-\sqrt{3}\)

OpenStudy (anonymous):

its −√3

OpenStudy (tkhunny):

Okay, then there is an angle such that its tangent is \(-\sqrt{3}\) For reference, there is an angle such that its tangent os \(\sqrt{3}\) I'd solve the second one, first. This will help you with the first one. You may wish to draw a Right Triangle and designate one fo the acute angles as your angle of interest.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks